# A Comprehensive Simulation Framework for CXL Disaggregated Memory

### v3

@@ V-D Future Work
Currently, CXL-DMSim only supports single-host simulation mode. Plans are on our schedule to enhance its capabilities to support multi-hosts via CXL switches with various CXL topologies, akin to dist-gem5 [54]. We are actively developing and integrating the CXL.cache sub-protocol into CXL-DMSim. Furthermore, we will leverage CXL-DMSim to explore heterogeneous memory management; e.g., efficient allocation and release of memory capacity and bandwidth within a CXL memory pool with tied and hybrid memory resources. These days, many researchers claim that the deployment of CXL-disaggregated memory in data centers offers both performance improvements and cost reductions. But, forecasting TCO savings and calculating Return on Investment (ROI) for the CXL technology are complex. We believe that leveraging CXL-DMSim for cost modeling, system efficiency evaluation, and optimization in a multi-host CXL-disaggregated memory environment are all interesting research topics.
@@ VI Conclusion
In this paper, we have presented CXL-DMSim, an open-source and silicon-calibrated full-system simulator for CXL disaggregated memory systems. CXL-DMSim can be used either in a NUMA-compatible kernel-managed mode or in an app-managed mode depending on users’ preference. Our experimental results on both hardware and CXL-DMSim suggest that CXL memory is very effective in memory expansion of both capacity and bandwidth to boost system performance. We also found the current CXL memory controller design composed of a CXL protocol controller and a DDR controller compromises both the latency and bandwidth of CXL memory and also leads to a high sensitivity to Rd/Wr patterns. This calls for a CXL-oriented memory controller design without DDR to fully unleash the performance of CXL links. With CXL-DMSim, the system performance bottlenecks can be quickly identified thanks to its high observability with substantial simulation statistics. In the future, we will continue to enhance the capabilities of CXL-DMSim, and we also welcome the community to join us to build a solid simulation platform for architectural research on memory-pooled computing systems.
Appendix A Artifact Appendix
A-A Abstract
The artifact consists of three components: 1) the complete CXL-DMSim source code based on gem5 v23.1; 2) a Linux kernel with the added CXL expander driver; 3) a disk image containing test programs. To initiate the full-system simulation of CXL-DMSim, the Linux kernel and disk image need to be loaded. This appendix describes the process of compiling and installing our simulator, CXL-DMSim, and reproducing the results shown in Fig. 10 and Fig. 15 in this paper. The code must be executed on a Linux system with at least 64GB of main memory and 50GB of disk space.
A-B Artifact Checklist (meta-information)
Program: We provide a disk image parsec.img that includes benchmarks LMBench and DLRM.
Compilation: GCC/G++ 9.4.0 and Python3.8.
Binary: The Linux 5.4.49 kernel binary with added CXL memory expender driver support.
Run-time environment: Ubuntu 20.04 or Ubuntu 22.04.
Hardware: Intel x86-64
Output: Files with the results of the program’s execution.
Experiments: Manual invocation of scripts, which launch corresponding experiments and generate outputs in designated folders.
How much disk space required (approximately)?: 40–50GB
How much time is needed to prepare workflow (approximately)?: The compilation for CXL-DMSim takes approximately 10-20 minutes, and the full-system bootup for CXL-DMSim takes about 30 minutes.
How much time is needed to complete experiments (approximately)?: We primarily control memory allocation on DDR-L, CXL memory, or interleave both using the numactl. For a single test of a specific memory allocation, LMBench requires 5 hours and Merci requires 20 hours. You can accelerate the testing progress by executing as many tests in parallel as possible, based on the size of the machine’s main memory.
Publicly available?: Yes
CXL-DMSim: https://github.com/ferry-hhh/CXL-DMSim.
Kernel and disk image: https://drive.google.com/drive/folders/1sxZBsedT19ntJdzN8MTkcbczMGXNXgrM?usp=sharing.
Code licenses (if publicly available)?: BSD 3-Clause “New” or “Revised” License
A-C Description
A-C1 How to access
All the source code of CXL-DMSim is available on GitHub, and the linux Kernel and disk image can be downloaded from GoogleDrive.
A-C2 Hardware dependencies
It is recommended to run this on a machine with a sufficiently large main memory (64GB) because modeling the operations of 48 O3 CPUs along with the CXL memory expender requires a considerable amount of memory.
A-C3 Software dependencies
Linux systems that support building gem5. Ubuntu 20.04 and Ubuntu 22.04 have been verified to run CXL-DMSim successfully. The GCC version used is 9.4.0, and the Python version is 3.8.10.
A-D Installation
1)
First, run the following command on Ubuntu 20.04 to install the required dependencies:
sudo apt install build-essential git m4 scons zlib1g zlib1g-dev libprotobuf-dev protobuf-compiler libprotoc-dev libgoogle-perftools-dev python3-dev python-is-python3 libboost-all-dev pkg-config
2)
Download the CXL-DMSim source code, linux kernel and disk image from the provided links.
3)
Run the following command to compile:
scons build/X86/gem5.opt -j{cpus}
Note that the ‘-j’ flag is optional and enables parallelization of the compilation process, with ‘cpus’ specifying the number of threads to use. It is recommended to use as many threads as possible to accelerate the compilation.
A-E Experiment Workflow
Next, we will proceed to complete the experiments for Fig. 10 and Fig. 15.
1)
Before starting, please check the correct paths for the kernel and disk_image within the file configs/example/gem5_library/x86-cxl-run.py on your local machine.
2)
At the terminal prompt, start the test script with the following command.
./run_fs.sh
The script will sequentially execute the LMBench and DLRM tests. However, you may also manually duplicate each test command within run_fs.sh to execute them in parallel, if the memory capacity of your host machine permits.
A-F Evaluation and Expected Results
Upon execution of the command to initiatethe the full system, the system runtime logs can be found within the board.pc.com_1.device file located in the corresponding test folder under the output directory. Subsequently, the experimental results will also be appended to this file. We provide the expected data for each test in the expected_result directory of the GitHub repository for your reference.
A-G Methodology
Submission, reviewing and badging methodology:

### v6

@@ V-D Future Work
CXL-DMSim now fully supports the CXL.io, CXL.cache, and CXL.mem sub-protocols, and multiple use cases for type-3 and type-1 devices have been implemented. Owing to space limitations, the calibration and implementation details of CXL.cache, along with related use cases (CXL-NIC), will be presented in another paper. Incremental support for type-2 use cases (CXL-GPU) is currently underway. Moreover, a single-host CXL switch has been realized to interconnect multiple CXL devices. Leveraging this foundation, we are progressing towards a distributed CXL-DMSim simulation platform, analogous to dist-gem5 [46]. In this framework, compute nodes, memory nodes, and switch-based fabric interconnect execute within distinct processes; packet exchanges and clock synchronization among these heterogeneous simulation nodes are facilitated through shared memory. Ultimately, this architecture enables a large-scale disaggregated memory system built upon a configurable CXL fabric.
@@ VI Conclusion
In this paper, we have presented CXL-DMSim, an open-source and silicon-calibrated full-system simulator for CXL disaggregated memory systems. CXL-DMSim can be used either in a NUMA-compatible kernel-managed mode or in an app-managed mode depending on users’ preference. Our experiments on both hardware and CXL-DMSim suggest that CXL memory is very effective in memory expansion of both capacity and bandwidth to boost system performance. With CXL-DMSim, the system performance bottlenecks can be quickly identified thanks to its high observability with substantial simulation statistics. In the future, we will continue to enhance the capabilities of CXL-DMSim, and we also welcome the community to join us in building a solid simulation platform for architectural research on memory-pooled computing systems.